OpenStudy (anonymous):

Anyone able to help me please! An explosion blows a rock into three pieces. Two pieces fly off at right angles to each other. Determine the mass and direction of the third piece. (drawing incoming).

OpenStudy (anonymous):



OpenStudy (anonymous):

so the Initial momentum =0

OpenStudy (anonymous):

After the explosion I got: PX] (1 kg * -12 m/s) + (3kg * 0 ) + ( m * Vx ??? ) I don't know what to do here

OpenStudy (anonymous):

PY] (1 kg * 0 ) + (3 kg * 8 m/s) + ( m * Vy ???? ) I don't know either

OpenStudy (anonymous):

that's all I could think of :(

OpenStudy (anonymous):

Px] -12 + m(Vx) = 0 Py] 24 + m(Vy) = 0

OpenStudy (anonymous):

12 + m * 30 m/s = 0 ???

OpenStudy (anonymous):

12 = -30m ???

OpenStudy (anonymous):

m = .4 kg ???

OpenStudy (anonymous):

i know it's wrong lol... ugh